OpenStudy (anonymous):

the velocity of a diver just before hitting the water is -9.3m/s, where the minus sign indicates tha her motion is directly downward. what is her displacement during the last 1.02s of the drive?

OpenStudy (anonymous):

The acceleration is -9.8 m/s^2. So, well, \[v = v _{0} + at\]\[v _{0} = -9.3 - (-9.8)(1.02) = 0.696 m/s\] Now, \[disp = v _{0}t + at ^{2}/2 = (0.696)(1.02) + (-9.8)(1.02)^{2}/2 = -11.486 m\] Your answer is -11.486 m